# Linear combinations of cluster mean-field states applied to spin systems

### v1

@@ Abstract
preprint: AIP/123-QED
I Introduction
Hartree-Fock (HF) mean-field theory treats complex systems using a wave function correct for non-interacting electrons. When the correlations between electrons are weak, this approach works well. When those correlations are strong, however, the HF treatment is inadequate. Cluster mean-field (cMF Jiménez-Hoyos and Scuseria (2015); Papastathopoulos-Katsaros et al. (2022); Papastathopoulos-Katsaros, Henderson, and Scuseria (2023); Ghassemi Tabrizi and Jiménez-Hoyos (2023)) theory generalizes this basic idea but provides a more nuanced and flexible framework. It also uses a wave function which is correct for non-interacting constituents, but where in HF these constituents are the individual electrons, cMF uses multi-electronic fragments. Crucially, the wave function within each fragment (which we usually refer to as a “cluster”, “tile”, “plaquette” or “covering”) is correlated. That is, cMF has a wave function which includes correlations within a fragment but not between fragments. This approach seems particularly suited to systems where correlations are in some sense localized, as seen in lattice models with short-range interactions or in systems of well-separated molecules.
In recent publications, Papastathopoulos-Katsaros et al. (2022); Papastathopoulos-Katsaros, Henderson, and Scuseria (2023) we have successfully applied cluster-based methodologies to address strongly correlated spin systems, demonstrating their effectiveness in dealing with the challenging $J_{1}-J_{2}$ and $XXZ$ Heisenberg models. These investigations are extensions of our previous work, where we introduced cluster mean-field theory (cMF) for fermionic systems.Jiménez-Hoyos and Scuseria (2015) One of the key questions we must ask in cMF is what the individual clusters should be. When the clusters are sufficiently large, the details of the cluster size and shape are not very important. However, because the computational cost of cMF scales exponentially with cluster size, practical considerations require relatively small clusters, and when the clusters are small, cMF is more sensitive to the size and shape of the tiles. We can try to mitigate this dependence to some extent by using inhomogeneous plaquettes,Neuscamman et al. (2011) but relying on a single underlying tiling scheme with relatively small tiles ultimately limits the accuracy of cMF.
@@ I Introduction
The basic idea of combining several tiling schemes is not a new one, but has mostly been restricted to two-site tiles (“dimers”) often associated with valence bond theory.Weyl, Rumer, and Teller (1932); Pauling (1989) This approach has found extensive applications in the examination of polymers and Heisenberg chains, Klein et al. (1997); Garcia-Bach (2000); Garcia-Bach, Valenti, and Klein (1997); Garcia-Bach and Klein (1996); Garcia-Bach, Peñaranda, and Klein (1992) and even square-planar lattices. Klein and Schmalz (1990); Klein et al. (1991) Notably, resonating valence bond theory Anderson (1987) has been put forward as a means to elucidate phenomena such as high-temperature superconductivity, particularly in cuprate compounds, rendering it even more relevant in the context of addressing strong-correlation effects.
A parallel endeavor was undertaken by Garcia-Bach and Klein, as evident in works like Refs. 13; 14, which are primarily concerned with singlet ground states and are limited to dimer coverings. In contrast, we expand our approach to accommodate diverse tile sizes and systems of varying $S^{2}$ quantum numbers. This is motivated as we want to explore both $XXZ$ and $J_{1}-J_{2}$ Heisenberg lattices, mainly because the intermediate spin-liquid-like region $J_{2}/J_{1}\approx 0.5$ of the square $J_{1}-J_{2}$ Heisenberg model is very difficult for conventional methods to describe and the $XXZ$ model does not have $S^{2}$ as a symmetry and comes with a diverse phase spectrum. For a more comprehensive understanding of the interplay between cMF and other advanced methodologies, as well as insights into the advantages of cMF relative to these approaches, readers are directed to Refs. 1; 2; 3 and the associated references therein.
Our focus here is on benchmarking cMF-based methods for spin systems because they have lower computational cost than do equivalently-sized fermionic lattices and because exact or nearly exact numerical results are readily available for relatively large spin lattices. It should, however, be noted that cMF has previously been shown to be effective for fermionic systems Jiménez-Hoyos and Scuseria (2015); Fang and Li (2007) as well as the spin systems which we are considering here. To assess the validity of our findings, we compare our results for one-dimensional (1D) systems to those obtained using the density matrix renormalization group (DMRG), Nakatani (2018) as it gives the exact results in 1D with a manageable, large bond dimension, and to those obtained using exact diagonalization (FCI) for two-dimensional systems.
@@
II Background
II.1 Heisenberg model
Spin lattices, particularly those represented by Heisenberg models, possess considerable chemical significance. An illustration of this is the analysis of iron-sulfur clusters, such as ferredoxins related to nitrogen fixation or photosynthesis, which have been simulated based on the Heisenberg model.Sharma et al. (2014) Another instance involves single-chain magnets, like Cobalt(II) Thiocyanate, which have been modeled using the $XXZ$ chain.Rams et al. (2020) Furthermore, specific electrides, conjugated hydrocarbons, and superconductors exhibit features reminiscent of Heisenberg exchange interactions.Wu, Schmalz, and Klein (2002); Dye (1997); Van Oosten, Broer, and Nieuwpoort (1995)
In this study, we investigate the $XXZ$ and the $J_{1}$ $J_{2}$ Heisenberg models. Both models characterize an assembly of interacting spins on a lattice with a finite size $N$. The $XXZ$ model specifically accounts for interactions solely between nearest neighbors and incorporates anisotropic interactions that violate $S^{2}$, whereas the $J_{1}$ $J_{2}$ model encompasses both nearest and next-nearest neighbor interactions with isotropic interactions, preserving $S^{2}$ as a symmetry. In the case of a one-dimensional lattice, both models can be exactly solved using Bethe ansatz.Bethe (1931); Yang and Yang (1966) Although exact solutions are not achievable for the 2D cases, extensive numerical studies have been conducted (refer to, for instance, Refs. 25; 26; 27; 28; 29; 30; 31; 32; 33; 34; 35; 36; 37; 38; 39; 40; 41; 42 for the $J_{1}-J_{2}$ Hamiltonian and Refs. 43; 44; 45; 46; 47; 48; 49 for the $XXZ$ model). The Hamiltonian of the $XXZ$ Heisenberg model is
@@ II.1 Heisenberg model
II.2 Cluster Mean-Field
@@
In this case, $\phi_{i}$ denotes cluster wavefunctions chosen to minimize $\bra{\Phi}H\ket{\Phi}$. It is worth noting that for SU(2) systems when the clusters consist of single sites, cMF simplifies to standard mean-field theory. However, for fermions, the optimization of the the single-site “tiles” is necessary for this to hold.Jiménez-Hoyos and Scuseria (2015)
@@
Apart from being influenced by the size and configuration of clusters, cMF also relies on the symmetry restrictions applied to these clusters. In this study, we maintain the constraint that each cluster represents an $S_{z}=0$ eigenstate (where we use the same symbol for the operator and the eigenvalues), consistent with the principles of both restricted (RcMF) and unrestricted cMF (UcMF). Despite our previous findings,Papastathopoulos-Katsaros, Henderson, and Scuseria (2023) which highlighted the nearly exact nature of generalized cMF (GcMF) around $\Delta=-1$ for the $XXZ$ model, our current emphasis lies in broadening our approach by incorporating additional cluster coverings rather than expanding the Hilbert space for each cluster.
@@
with the requirement of the $c^{ij}$ coefficient optimization. This increases the computational cost, but provides much larger variational freedom. In our work, we will use the former definition for dimers and the more general approach (with coefficient optimization) for tetramers, for which a complete basis cannot be straightforwardly defined as we shall see in section IV.2.
II.3 Matrix elements and cMF optimization
The computation of matrix elements in LC-cMF follows a similar procedure to what we have described in our prior publications.Papastathopoulos-Katsaros et al. (2022); Papastathopoulos-Katsaros, Henderson, and Scuseria (2023) Calculating matrix elements involving states with distinct tilings does not significantly increase the computational cost. The overall efficiency of this approach is contingent on the number of coverings included and the optimization strategy for the coefficients (whether by selecting specific singlet and triplet states or by concurrently optimizing the coefficients for each covering).
II.4 Computational details
In this study, our computations were conducted using an in-house code, which utilizes the ITensor Fishman, White, and Stoudenmire (2022) library to construct the cMF states, simplifying the computation of necessary matrix elements. In the tetramer scenario, we initiated the cluster states with random configurations and employed the conjugate gradient Atkinson (1989) algorithm with numerical gradients to optimize the energy. For conducting the NOCI, we utilized the generalized eigenvalue solver from the GNU Scientific Library (GSL).
@@ II.4 Computational details
III LC-cMF
III.1 A complete basis of dimer coverings
A “covering” or “tiling” is defined as a product of cluster states with non-overlapping indices. It has been shown Garcia-Bach, Valenti, and Klein (1997) that a linear combination of a combinatorial number of linearly independent dimer coverings of valence bonds (RcMF[2]) with optimized coefficients can give the exact singlet ground state of Heisenberg lattices. Mathematically,
@@ III.1 A complete basis of dimer coverings
where $N$ is the number of sites. Finding these dimer coverings is non-trivial and constitutes a problem in graph theory and statistical physics.Kasteleyn (1961) Below we will show the algorithm for the 1D case, and for the interested reader, the 2D case is extensively covered in Refs. 13; 14.
@@
Figure 3: All 5 dimer coverings for the 6-site Heisenberg chain. Figure adapted from Ref. 10.
III.2 Truncating the covering basis
@@
To address this issue, we opted for a strategy involving the selection of tilings characterized by relatively short “bond lengths”, where “bond length” represents the summation of distances between sites within each dimer for all the dimers in a given covering. Thus, a nearest-neighbor dimer contributes a bond length of 1, a third-nearest-neighbor dimer contributes a bond length of 3, and so forth; recall that there are no next-nearest-neighbor dimers by virtue of the bipartition of the lattice.
Once we have established the various tilings, we can order them by their bond lengths. Tab. 1 shows the results of this process for the dimer coverings in Fig. 3. We can then define a “bond level” which denotes the position of the bond length within this ordering. In the context of Fig. 3, bond level 1 refers to tiling |$\ket{s_{1}}$, bond level 2 refers to tilings |$\ket{s_{2}}$ and |$\ket{s_{3}}$, and so on. We can now define a truncation wherein we retain all tilings within a certain bond level. We symbolize this truncation scheme as $d$LC-cMF[$n$], where $d$ refers to the bond level and $n$ to the cluster size. For example, 3LC-cMF[2] refers to a computation utilizing all dimer coverings ($n=2$) with a bond level of 3 or less ($d=3$).
For 1D chains with open boundary conditions (OBC), the number of dimer coverings with bond level of 1 scales as $\mathcal{O}(N)$ while for bond level of 2 it scales as $\mathcal{O}(N^{2})$. For lattices with periodic boundary conditions (PBC) and 2D lattices, computing the bond length needs to take the periodicity into account. For example, in PBC, the first and the last site have only 1 site distance, the second to last with the last only 2, etc., whereas, in 2D lattices, the length can be computed by using the smallest bond length while moving along each axis of the lattice. Again, the bond level is defined similarly, regardless of the actual value of the bond lengths.
@@
To demonstrate the effectiveness of the dimer covering basis, we begin our analysis by computing the ground state energy of a 16-site $J_{1}-J_{2}$ Heisenberg chain with OBC. The results are presented in Fig. 4. We remind the reader that 2-LC means we include the nearest neighbor tiling and the $N$ tilings with equivalent next shortest bond lengths overall. As mentioned in Sec. II, the exact ground state at the Majumdar-Ghosh point ($J_{2}/J_{1}=0.5$) is a nearest-neighbor dimer covering, hence our method also ought to be exact. However, it is also evident that even with a minimal bond level, our results exhibit qualitatively good agreement with the exact ground state energy, particularly in the vicinity of $J_{2}/J_{1}=0.5$. Moreover, the improvement in comparison to a UcMF[2] calculation, where the coefficients are optimized, is significant, underscoring the considerable potential of LC-cMF, which comes without the need of any additional “correlation” methods; the correlations are described by a simple linear combination of tilings. Finally, it is worth noting that while UcMF[2] has broken $S^{2}$ symmetry, 2LC-cMF[2] is an $S^{2}$ eigenstate. Consequently, this energy improvement comes with the correct quantum numbers, in contrast to UcMF[2].
@@
To illustrate that dimer coverings can also be effectively applied to 2D systems, we take the example of the $4\times 4$ $J_{1}-J_{2}$ Heisenberg square with PBC, and our findings are presented in Fig. 5. An intriguing observation is that, once again, we achieve remarkable accuracy for such a simple wave function, especially when compared to UcMF[$2\times 2$], which has the flexibility of coefficient optimization. This underscores that, with relatively small clusters, LC-cMF can outperform the results obtained with larger clusters. In particular, in the non-magnetic $J_{2}/J_{1}\approx 0.5$ region, LC-cMF[2] excels, which is the region that is most difficult to describe. This is probably a spin-liquid-like feature that this regime has, which can be captured by this ansatz.
@@
IV Correlation and non-singlet cases

### v2

@@ Abstract
1 Introduction
Hartree-Fock (HF) mean-field theory treats complex systems using a wave function correct for non-interacting electrons. When the correlations between electrons are weak, this approach works well. When those correlations are strong, however, the HF treatment is inadequate. Cluster mean-field (cMF 1, 2, 3, 4, 5, 6, 7, 8, 9) theory generalizes this basic idea but provides a more nuanced and flexible framework. It also uses a wave function which is correct for non-interacting constituents, but where in HF these constituents are the individual electrons, cMF uses multi-electronic fragments. Crucially, the wave function within each fragment (which we usually refer to as a “cluster”, “tile”, “plaquette” or “covering”) is correlated. That is, cMF has a wave function which includes correlations within a fragment but not between fragments. This approach seems particularly suited to systems where correlations are in some sense localized, as seen in lattice models with short-range interactions or in systems of well-separated molecules.
In recent publications, 2, 3 we have successfully applied cluster-based methodologies to address strongly correlated spin systems, demonstrating their effectiveness in dealing with the challenging $J_{1}-J_{2}$ and $XXZ$ Heisenberg models. These investigations are extensions of our previous work, where we introduced cluster mean-field theory (cMF) for fermionic systems.1 One of the key questions we must ask in cMF is what the individual clusters should be. When the clusters are sufficiently large, the details of the cluster size and shape are not very important. However, because the computational cost of cMF scales exponentially with cluster size, practical considerations require relatively small clusters, and when the clusters are small, cMF is more sensitive to the size and shape of the tiles. We can try to mitigate this dependence to some extent by using inhomogeneous plaquettes,10 but relying on a single underlying tiling scheme with relatively small tiles ultimately limits the accuracy of cMF.
@@ 1 Introduction
The basic idea of combining several tiling schemes is not a new one, but has mostly been restricted to two-site tiles (“dimers”) often associated with valence bond theory.11, 12 This approach has found extensive applications in the examination of polymers and Heisenberg chains, 13, 14, 15, 16, 17 and even square-planar lattices. 18, 19 Notably, resonating valence bond theory 20 has been put forward as a means to elucidate phenomena such as high-temperature superconductivity, particularly in cuprate compounds, rendering it even more relevant in the context of addressing strong-correlation effects.
A parallel endeavor was undertaken by Garcia-Bach and Klein, as evident in works like Refs. 18, 19, which are primarily concerned with singlet ground states and are limited to dimer coverings. In contrast, we expand our approach to accommodate diverse tile sizes and systems of varying $S^{2}$ quantum numbers. This is motivated as we want to explore both $XXZ$ and $J_{1}-J_{2}$ Heisenberg lattices, mainly because the intermediate spin-liquid-like region $J_{2}/J_{1}\approx 0.5$ of the square $J_{1}-J_{2}$ Heisenberg model is very difficult for conventional methods to describe and the $XXZ$ model does not have $S^{2}$ as a symmetry and comes with a diverse phase spectrum. For a more comprehensive understanding of the interplay between cMF and other advanced methodologies, as well as insights into the advantages of cMF relative to these approaches, readers are directed to Refs. 1, 2, 3 and the associated references therein.
Our focus here is on benchmarking cMF-based methods for spin systems because they have lower computational cost than do equivalently-sized fermionic lattices and because exact or nearly exact numerical results are readily available for relatively large spin lattices. It should, however, be noted that cMF has previously been shown to be effective for fermionic systems 1, 21 as well as the spin systems which we are considering here. To assess the validity of our findings, we compare our results for one-dimensional (1D) systems to those obtained using the density matrix renormalization group (DMRG), 22 as it gives the exact results in 1D with a manageable, large bond dimension, and to those obtained using exact diagonalization (FCI) for two-dimensional systems.
@@
2 Background
2.1 Heisenberg model
Spin lattices, particularly those represented by Heisenberg models, possess considerable chemical significance. An illustration of this is the analysis of iron-sulfur clusters, such as ferredoxins related to nitrogen fixation or photosynthesis, which have been simulated based on the Heisenberg model.23 Another instance involves single-chain magnets, like Cobalt(II) Thiocyanate, which have been modeled using the $XXZ$ chain.24 Furthermore, specific electrides, conjugated hydrocarbons, and superconductors exhibit features reminiscent of Heisenberg exchange interactions.25, 26, 27
In this study, we investigate the $XXZ$ and the $J_{1}$ $J_{2}$ Heisenberg models. Both models characterize an assembly of interacting spins on a lattice with a finite size $N$. The $XXZ$ model specifically accounts for interactions solely between nearest neighbors and incorporates anisotropic interactions that violate $S^{2}$, whereas the $J_{1}$ $J_{2}$ model encompasses both nearest and next-nearest neighbor interactions with isotropic interactions, preserving $S^{2}$ as a symmetry. In the case of a one-dimensional lattice, both models can be exactly solved using Bethe ansatz.28, 29 Although exact solutions are not achievable for the 2D cases, extensive numerical studies have been conducted (refer to, for instance, Refs. 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47 for the $J_{1}-J_{2}$ Hamiltonian and Refs. 48, 49, 50, 51, 52, 53, 54 for the $XXZ$ model). The Hamiltonian of the $XXZ$ Heisenberg model is
@@ 2.1 Heisenberg model
2.2 Cluster Mean-Field
@@
In this case, $\phi_{i}$ denotes cluster wavefunctions chosen to minimize $\bra{\Phi}H\ket{\Phi}$. It is worth noting that for SU(2) systems when the clusters consist of single sites, cMF simplifies to standard mean-field theory. However, for fermions, the optimization of the the single-site “tiles” is necessary for this to hold.1
@@
Apart from being influenced by the size and configuration of clusters, cMF also relies on the symmetry restrictions applied to these clusters. In this study, we maintain the constraint that each cluster represents an $S_{z}=0$ eigenstate (where we use the same symbol for the operator and the eigenvalues), consistent with the principles of both restricted (RcMF) and unrestricted cMF (UcMF). Despite our previous findings,3 which highlighted the nearly exact nature of generalized cMF (GcMF) around $\Delta=-1$ for the $XXZ$ model, our current emphasis lies in broadening our approach by incorporating additional cluster coverings rather than expanding the Hilbert space for each cluster.
@@
with the requirement of the $c^{ij}$ coefficient optimization. This increases the computational cost, but provides much larger variational freedom. In our work, we will use the former definition for dimers and the more general approach (with coefficient optimization) for tetramers, for which a complete basis cannot be straightforwardly defined as we shall see in section 4.2.
2.3 Matrix elements and cMF optimization
The computation of matrix elements in LC-cMF follows a similar procedure to what we have described in our prior publications.2, 3 Calculating matrix elements involving states with distinct tilings does not significantly increase the computational cost. The overall efficiency of this approach is contingent on the number of coverings included and the optimization strategy for the coefficients (whether by selecting specific singlet and triplet states or by concurrently optimizing the coefficients for each covering).
2.4 Computational details
In this study, our computations were conducted using an in-house code, which utilizes the ITensor 67 library to construct the cMF states, simplifying the computation of necessary matrix elements. In the tetramer scenario, we initiated the cluster states with random configurations and employed the conjugate gradient 68 algorithm with numerical gradients to optimize the energy. For conducting the NOCI, we utilized the generalized eigenvalue solver from the GNU Scientific Library (GSL).
@@ 2.4 Computational details
3 LC-cMF
3.1 A complete basis of dimer coverings
A “covering” or “tiling” is defined as a product of cluster states with non-overlapping indices. It has been shown 15 that a linear combination of a combinatorial number of linearly independent dimer coverings of valence bonds (RcMF[2]) with optimized coefficients can give the exact singlet ground state of Heisenberg lattices. Mathematically,
@@ 3.1 A complete basis of dimer coverings
where $N$ is the number of sites. Finding these dimer coverings is non-trivial and constitutes a problem in graph theory and statistical physics.69 Below we will show the algorithm for the 1D case, and for the interested reader, the 2D case is extensively covered in Refs. 18, 19.
@@
Figure 3: All 5 dimer coverings for the 6-site Heisenberg chain. Figure adapted from Ref. 15.
3.2 Truncating the covering basis
@@
To address this issue, we begin by noting that the Hamiltonian only directly couples nearest-neighbor and next-nearest-neighbor sites. We therefore expect the most significant correlations to be relatively short range, suggesting that the most important tilings should be fairly localized. For this reason, we opted for a strategy involving the selection of tilings characterized by relatively short “bond lengths”, where “bond length” represents the summation of distances between sites within each dimer for all the dimers in a given covering. Thus, a nearest-neighbor dimer contributes a bond length of 1, a third-nearest-neighbor dimer contributes a bond length of 3, and so forth (recall that there are no next-nearest-neighbor dimers by virtue of the bipartition of the lattice).
Once we have established the various tilings, we can order them by their bond lengths. Table 1 shows the results of this process for the dimer coverings in Fig. 3. We can then define a “bond level” which denotes the position of the bond length within this ordering. In the context of Fig. 3, bond level 1 refers to tiling |$\ket{s_{1}}$, bond level 2 refers to tilings |$\ket{s_{2}}$ and |$\ket{s_{3}}$, and so on. We can now define a truncation wherein we retain all tilings within a certain bond level. Increasing the bond level increases the number of tilings we include, and we always include the most local tilings first. We symbolize this truncation scheme as $d$LC-cMF[$n$], where $d$ refers to the bond level and $n$ to the cluster size. For example, 3LC-cMF[2] refers to a computation utilizing all dimer coverings ($n=2$) with a bond level of 3 or less ($d=3$).
For 1D chains with open boundary conditions (OBC), the number of dimer coverings with bond level of 1 scales as $\mathcal{O}(N)$ while for bond level of 2 it scales as $\mathcal{O}(N^{2})$. For lattices with periodic boundary conditions (PBC) and 2D lattices, computing the bond length needs to take the periodicity into account. For example, in PBC, the first and the last site have only 1 site distance, the second to last with the first only 2, etc., whereas, in 2D lattices, the length can be computed by using the smallest bond length while moving along each axis of the lattice. Again, the bond level is defined similarly, regardless of the actual value of the bond lengths.
Finally, we should note where cMF is extensive (the energy scales linearly with system size), LC-cMF as we have employed it here is not. While extensivity is an important consideration for calculations which seek to reach the thermodynamic limit, it is less important for the relatively small systems we have considered in this work. Generalizations of the schemes we have introduced here to reduce or eliminate extensivity errors are of course highly desirable and are under investigation.
@@
To demonstrate the effectiveness of the dimer covering basis, we begin our analysis by computing the ground state energy of a 16-site $J_{1}-J_{2}$ Heisenberg chain with OBC. The results are presented in Fig. 4. We remind the reader that 2-LC means we include the nearest neighbor tiling and the $N$ tilings with equivalent next shortest bond lengths overall. As mentioned in Sec. 2, the exact ground state at the Majumdar-Ghosh point ($J_{2}/J_{1}=0.5$) is a nearest-neighbor dimer covering, hence our method also ought to be exact. However, it is also evident that even with a minimal bond level, our results exhibit qualitatively good agreement with the exact ground state energy, particularly in the vicinity of $J_{2}/J_{1}=0.5$. Moreover, the improvement in comparison to a UcMF[2] calculation, where the coefficients are optimized, is significant, underscoring the considerable potential of LC-cMF, which comes without the need of any additional “correlation” methods; the correlations are described by a simple linear combination of tilings. Finally, it is worth noting that while UcMF[2] has broken $S^{2}$ symmetry, 2LC-cMF[2] is an $S^{2}$ eigenstate. Consequently, this energy improvement comes with the correct quantum numbers, in contrast to UcMF[2].
@@
To illustrate that dimer coverings can also be effectively applied to 2D systems, we take the example of the $4\times 4$ $J_{1}-J_{2}$ Heisenberg square with PBC, and our findings are presented in Fig. 5. An intriguing observation is that, once again, we achieve remarkable accuracy for such a simple wave function, especially when compared to UcMF[$2\times 2$], which has the flexibility of coefficient optimization. This underscores that, with relatively small clusters, LC-cMF can outperform the results obtained with larger clusters. In particular, LC-cMF[2] excels in the non-magnetic $J_{2}/J_{1}\approx 0.5$ region, which is the region that is most difficult to describe. This is probably a spin-liquid-like feature that this regime has, which can be captured by this ansatz.
@@
4 Correlation and non-singlet cases